OpenStudy (anonymous):

can someone PLEASE give me the formula for this kind of problem??! find the first five terms of the given sequence. a1=(4/5), a(sub)n+1=((n+1)/(n^2))*a(sub)n

OpenStudy (btaylor):

basically, plug in different values of n. so, for n=1, \[a_{1+1}=a_2 = \frac{2}{1} \times \frac{4}{5} = \frac{8}{5} = 1.6\]Now do the same thing, with n=2 to get n=3, etc.